OpenStudy (wwe123):

in an arithmetic progression, the sum of p terms is m and the sum of q terms is also m.Find the sum of (p+q) terms

hartnn (hartnn):

know the formula for sum of 'n' terms ?

OpenStudy (wwe123):

yes

hartnn (hartnn):

did u try to equate sum of p and sum of q terms...

OpenStudy (wwe123):

yes

hartnn (hartnn):

did the factor of p-q cancel out ?

hartnn (hartnn):

and is the answer = 0 ??

OpenStudy (wwe123):

answer is 0 ,but how to show ,i am not able to do...

hartnn (hartnn):

p[2a+(p-1)d]= q[2a+(q-1)d] did u get this ?

hartnn (hartnn):

this u get if u know the formula.

OpenStudy (wwe123):

yes, i know the formua but (p+q) ??

hartnn (hartnn):

did u expand that ?

hartnn (hartnn):

what u got ?

OpenStudy (wwe123):

1 min

hartnn (hartnn):

there is a factor of p-q which gets cancelled.

OpenStudy (wwe123):

how to expand and solve @hartnn

hartnn (hartnn):

p[2a+(p-1)d]= q[2a+(q-1)d] 2ap + p(p-1)d = 2aq + q(q-1)d 2a(p-q) = d[q(q-1)-p(p-1)] did u get this? want to try further ?

hartnn (hartnn):

u get p-q on right side also.

hartnn (hartnn):

did u get or should i continue ?

OpenStudy (wwe123):

plz continue, i am stunk

hartnn (hartnn):

its simple algebra..... 2a(p-q) = d[q(q-1)-p(p-1)] 2a(p-q) = d[q^2-q-p^2+p] 2a(p-q) = d[q^2-p^2 -(q-p] 2a(p-q) = d(q-p)[q+p-1] 2a(p-q) =- d(p-q)[q+p-1] we can cancel p-q because p-q is not 0, because p not =q 2a + (p+q-1)d = 0

hartnn (hartnn):

i hope u can do the last step....

OpenStudy (wwe123):

NOW sum upto (p+q) terms = (p+q)/2(2a+(p+q-1)d) =(p+q)/2(2a-2a) =0 is this right @hartnn

hartnn (hartnn):

yes. good :)

hartnn (hartnn):

clear with all steps ?

OpenStudy (wwe123):

thanks :)

hartnn (hartnn):

welcome ^_^